# DeepEye-SQL: A Software-Engineering-Inspired Text-to-SQL Framework

### v2

@@ 4. The Design Details of DeepEye-SQL
In this section, we provide a detailed exposition of the technical implementation of each phase within the DeepEye-SQL framework. We will describe the specific algorithms and design rationales that underlie our methodology.
4.1. Semantic Value Retrieval
Value Embedding. For each selected column $C_{j}$, we extract its unique values. Each distinct value $v$ is then encoded into a high-dimensional vector representation $\mathbf{e}_{v}$ using a pretrained sentence embedding model, specifically Qwen3-Embedding-0.6B (Zhang et al., 2025b). This embedding transforms discrete text strings into continuous semantic vectors, where values with similar meanings are located closer to each other in the vector space.
Vector Indexing. To facilitate fast similarity search, the generated value embeddings for each column are used to build a vector index. We employ Chroma (chroma core, 2025) with the Hierarchical Navigable Small World (HNSW) algorithm (Malkov and Yashunin, 2020) for this purpose. HNSW is highly efficient for approximate nearest neighbor (ANN) search, making it ideal for real-time applications. The result is a set of persistent, per-column vector indices $\{\mathcal{I}_{1},\mathcal{I}_{2},...,\mathcal{I}_{M}\}$, where each $\mathcal{I}_{j}$ is an indexed collection of all value embeddings for a column $C_{j}$.
Online Retrieval.
@@ 4.1. Semantic Value Retrieval
During the online phase, when a user question $Q$ is received, the system retrieves the most relevant values for that question from the pre-built indices. This process is detailed in Algorithm 1.
@@ 4.2. Robust Schema Linking
Schema linking, the task of identifying the correct subset of tables and columns relevant to a user’s question, is a cornerstone of any Text-to-SQL system (Liu et al., 2025a). Existing methods (Li et al., 2024b, 2025b; Talaei et al., 2024) often treat this as a direct mapping task, which can be brittle when faced with complex schemas or ambiguous questions. An error at this stage is catastrophic, as an incomplete or incorrect schema makes generating a correct SQL query nearly impossible. This critical dependency is analogous to the role of a formal specification (Hierons et al., 2009) in software development; without an accurate specification, the final product is destined to fail. To address this, we introduce a Robust Schema Linking module that, inspired by the principle of fault tolerance, combines multiple, diverse strategies to ensure the most accurate and complete schema is identified. Our overall process is detailed in Algorithm 2.
@@
However, $\mathcal{D}_{union}$ may be an incomplete, disconnected graph of schema elements. To resolve this, we enforce relational closure. We parse all foreign key relationships in the database schema $\mathcal{D}$. For any pair of tables $(T_{i},T_{j})$ present in $\mathcal{D}_{union}$, we automatically add the corresponding primary and foreign key columns that link them. This step guarantees that the final linked schema is a connected graph, providing a solid foundation for constructing joins in the SQL generation phase. This final, closed schema is defined as:
@@ 4.3. N-version Programming for SQL Generation
ICL-based SQL Generation. This generator leverages in-context learning (ICL), analogous to case-based reasoning (Kolodner, 1993). By providing the LLM with relevant examples, we ground its generation in proven patterns. The process involves retrieving schema-masked similar examples from a training set following DAIL-SQL (Gao et al., 2024), instructing the LLM to identify a common pattern, and then applying that pattern to the target question. This is formalized as:
@@ 4.4. SQL Unit Testing and Revision via Tool-Chain
SQL queries generated by LLMs, while often structurally sound, can contain minor yet critical errors (Liu et al., 2025b), such as incorrect function usage or logical flaws in JOIN conditions. A well-known limitation of LLMs is their unreliability in self-correction (Qu et al., 2025); when asked to review their own output, they exhibit a strong confirmation bias and tend to overlook their mistakes (Zhang et al., 2025c). This challenge is directly analogous to a core principle in software engineering: code should be validated by an independent testing suite, not just by the developer who wrote it. To address this, our framework introduces the SQL Unit Testing and Revision via Tool-Chain phase, which draws direct inspiration from the practice of Unit Testing (Runeson, 2006). In this paradigm, individual software components—or “units”—are tested in isolation. We adapt this concept by treating each functional component of a SQL query (e.g., its JOIN logic, a specific clause, or its syntax) as a testable “unit”. The Tool-Chain operationalizes this principle by externalizing the verification process. It uses a deterministic chain of specialized checkers, where each checker acts as an automated test case for a specific SQL unit. This allows the system to systematically detect errors and guide the LLM through a reliable, targeted revision process, as detailed in Algorithm 4.
@@
Table 1. Specialized Checkers.
@@
Syntax Checker SQL syntax and SELECT * FORM table (typo)
execution errors WHERE column = (incomplete)
JOIN Checker Non-standard ON cond1 OR cond2
JOIN conditions ON column IN (...)
ORDER-BY Checker Invalid ORDER BY ORDER BY MAX(...) LIMIT 1
+ LIMIT combinations ORDER BY COUNT(*) LIMIT 3
Time Checker Incorrect time function STRFTIME(’%Y’, date) format
usage and formatting Invalid date comparisons
SELECT Checker Ambiguous return SELECT * → specific columns
columns Unnecessary wildcards
MaxMin Checker Suboptimal MAX/MIN WHERE col = (SELECT MAX(...))
patterns ORDER BY col DESC LIMIT 1
NULL Checker NULL values in Add WHERE col IS NOT NULL
ORDER BY columns for ORDER BY columns
Result Checker Empty or meaningless Queries returning only NULL
results Zero-row result sets
Our approach centers on a suite of specialized, deterministic programs we term “Checker Tools”. These checkers are not applied randomly; they are organized into a sequential tool-chain that processes each SQL query in a specific order to catch errors systematically, from fundamental syntax to semantic correctness. The complete, ordered sequence of checkers and their functions is detailed in Table 1. The chain begins with the most fundamental validation: the Syntax Checker ensures the query is syntactically valid and executable, catching basic typos or incomplete clauses. It then proceeds to clause-specific validation. For example, the JOIN Checker flags non-standard conditions, the ORDER-BY Checker corrects invalid patterns like using aggregations with LIMIT, and the Time Checker validates date and time formatting. Subsequently, the chain addresses more semantic and stylistic issues. The SELECT Checker replaces ambiguous wildcards like SELECT * with specific column names, and the MaxMin Checker refactors suboptimal patterns into more efficient ORDER BY ... LIMIT 1 clauses. Finally, the NULL Checker and Result Checker add guards for potential NULL issues and flag queries that are likely to produce empty or meaningless results, ensuring the final output is not only correct but also useful.
@@
    The LLM is then invoked in a special “revision mode”. It receives the original context, the faulty SQL $\mathcal{S}_{cand}$, and the explicit directive $d_{err}$ from the checker. The LLM’s task is not to find the error, but to fix it based on the directive. This can be formalized as:
@@ 4.5. Confidence-aware SQL Selection
The preceding phases of our framework produce a set of high-quality, revised SQL candidates, $\mathcal{C}_{revised}$. However, these candidates may not be identical and could yield different execution results. The most common approach for selecting a final query is self-consistency (Xie et al., 2025; Li et al., 2025b; Gao et al., 2024), where all candidates are executed, and the query corresponding to the most frequent result is chosen. This method, while a strong baseline, has a critical flaw: the most popular answer is not always the correct one, especially for complex problems where multiple generation paths might converge on the same plausible but incorrect logic. This final challenge mirrors the release stage in a software lifecycle, which is governed by a Quality Gate (Schnell and Reinhart, 2016). A quality gate’s purpose is not merely to accept the most-voted-for version, but to enforce a set of objective quality criteria before a product is released. Similarly, our Confidence-aware SQL Selection phase acts as this quality gate for the generated SQL. It overcomes the flaws of simple majority voting by using the initial vote’s confidence as a quality metric to guide a more reliable, adaptive selection process.
@@
(1) Establish Unbalanced Cognitive Prior. First, instead of a neutral pairwise comparison, the LLM is explicitly primed with a “cognitive prior”. It is informed that, based on execution results, $\mathcal{S}_{i}$ has a higher prior confidence than $\mathcal{S}_{j}$. We instructs the LLM to select the higher-confidence candidate unless there is clear and compelling evidence that it is incorrect or that the lower-confidence candidate is demonstrably superior.
(2) Perform Pairwise Adjudication. Next, the LLM performs pairwise comparisons for the top-$K$ candidates. To ensure reliability, we employ a self-consistency mechanism. For each pair $(\mathcal{S}_{i},\mathcal{S}_{j})$, we sample multiple judgments and define the final stable vote, $V(\mathcal{S}_{i},\mathcal{S}_{j})$, as the majority outcome. The vote $V(\mathcal{S}_{i},\mathcal{S}_{j})$ yields 1 if $\mathcal{S}_{i}$ is superior and 0 if $\mathcal{S}_{j}$ is superior. From these consistent pairwise results, we compute an aggregate win rate for each candidate $\mathcal{S}_{i}$ as its average score against all other competitors in the top-$K$ set:
@@ 5.1. Experimental Setup
Datasets. We evaluate the performance of DeepEye-SQL on two widely-recognized and challenging cross-domain Text-to-SQL benchmarks: BIRD (Li et al., 2024a) and Spider (Yu et al., 2018). BIRD is a large-scale benchmark designed to mirror complex, real-world scenarios. It contains 12,751 unique question-SQL pairs across 95 databases from over 37 professional domains. Its databases are notably large and feature messy data and intricate schemas, making it a difficult test for grounding and robustness. Spider is a foundational large-scale, cross-domain benchmark in the field. It consists of 10,181 questions and 5,693 unique, complex SQL queries across 200 databases covering 138 different domains. Following prior works (Pourreza et al., 2025; Liu et al., 2025c; Cao et al., 2024), we use the development set of BIRD (BIRD-Dev) and the test set of Spider (Spider-Test) for our main evaluation.
@@
Evaluation Metrics. Following prior works (Pourreza et al., 2025; Liu et al., 2025c; Cao et al., 2024), our primary evaluation metric is Execution Accuracy (EX). A generated SQL query is considered correct if its execution result is equivalent to that of the ground-truth query after accounting for ordering. To measure the potential of our N-version Programming for SQL Generation module, we report the Upper-bound EX, which is the execution accuracy an oracle would achieve by always selecting the correct SQL from the generated candidates (Pourreza et al., 2025). For evaluating the Robust Schema Linking module, we use Table/Column Recall, the proportion of ground-truth tables and columns correctly identified. Finally, to assess practical efficiency, we measure Token Cost, corresponding to the number of tokens processed by LLMs.
@@
Baselines. We compare DeepEye-SQL against SOTA baselines from two paradigms (Table 2). The first is fine-tuning-based methods, which are trained on in-domain data, including strong competitors like XiYan-SQL (Liu et al., 2025c), CHASE-SQL (Pourreza et al., 2025), and OmniSQL (Li et al., 2025a). The second is prompting-based methods, which, like DeepEye-SQL, are training-free. This group includes methods like Alpha-SQL (Li et al., 2025b), RSL-SQL (Cao et al., 2024), and CHESS (Talaei et al., 2024).
@@
Implementation Details. All experiments are conducted on an Ubuntu 22.04.3 LTS server with 512GB of RAM and dual 40-core Intel(R) Xeon(R) Platinum 8383C CPUs. We deploy all open-source LLMs locally using the vllm (Kwon et al., 2023) framework on 4 NVIDIA A100 GPUs, each with 80GB of memory, and accelerate inference using a tensor parallelism of 4. To validate the robustness and generalizability of our framework, we integrated DeepEye-SQL with three distinct models from two leading series: Gemma3-27B-Instruct (Kamath et al., 2025), Qwen2.5-Coder-32B-Instruct (Hui et al., 2024), and Qwen3-Coder-30B-A3B-Instruct (Yang et al., 2025). Unless otherwise specified, the pipeline is configured as follows. For Semantic Value Retrieval, where we retrieve up to the top-5 most similar values for each TEXT-type column, we use the Qwen3-Embedding-0.6B (Zhang et al., 2025b) model, and the similarity threshold ($\theta_{val}$) for Value-based Schema Linking is set to 0.98. The draft SQL for Reversed Schema Linking is produced by the ICL-based generator. For each LLM sub-task (e.g., Direct Schema Linking), we set the sampling budget to 8 with a sampling temperature of 0.7. Finally, in the Confidence-aware SQL Selection phase, which adjudicates between the top-2 ranked queries in low-confidence scenarios, the confidence shortcut threshold ($\theta_{conf}$) is set to 0.6. All prompt templates of our method are in the Appendix.

### v3

@@ 3. The Design Details of DeepEye-SQL
3.1. Semantic Value Retrieval
Value Embedding. For each selected column $C_{j}$, we extract its unique values. Each distinct value $v$ is then encoded into a high-dimensional vector representation $\mathbf{e}_{v}$ using a pretrained sentence embedding model, specifically Qwen3-Embedding-0.6B (Zhang et al., 2025c). This embedding transforms discrete text strings into continuous semantic vectors, where values with similar meanings are located closer to each other in the vector space.
Vector Indexing. To facilitate fast similarity search, the generated value embeddings for each column are used to build a vector index. We employ Chroma (chroma-core, 2025) with the Hierarchical Navigable Small World (HNSW) algorithm (Malkov and Yashunin, 2020) for this purpose. HNSW is highly efficient for approximate nearest neighbor (ANN) search, making it ideal for real-time applications. The result is a set of persistent, per-column vector indices $\{\mathcal{I}_{1},\mathcal{I}_{2},...,\mathcal{I}_{M}\}$, where each $\mathcal{I}_{j}$ is an indexed collection of all value embeddings for a column $C_{j}$.
Online Retrieval. During the online phase, when a user question $Q$ is received, the system retrieves the most relevant values for that question from the pre-built indices. This process is detailed in Algorithm 1.
@@ 3.2. Robust Schema Linking
Schema linking, the task of identifying the correct subset of tables and columns relevant to a user’s question, is a cornerstone of any Text-to-SQL system (Liu et al., 2025a). Existing methods (Li et al., 2024a, 2025a; Talaei et al., 2024) often treat this as a direct mapping task, which can be brittle when faced with complex schemas or ambiguous questions. An error at this stage is catastrophic, as an incomplete or incorrect schema makes generating a correct SQL query nearly impossible. This critical dependency is analogous to the role of a formal specification (Hierons et al., 2009) in software development; without an accurate specification, the final product is destined to fail. To address this, we introduce a Robust Schema Linking module that, inspired by the principle of fault tolerance, combines multiple, diverse strategies to ensure the most accurate and complete schema is identified. Our overall process is detailed in Algorithm 2.
@@
However, since $\mathcal{D}_{union}$ is aggregated from independent strategies, it acts effectively as a collection of isolated schema nodes (tables). Without explicit edges (Foreign Keys) to connect them, the resulting schema graph is often disconnected, rendering valid JOIN operations impossible. To resolve this, we enforce relational closure as a mandatory connectivity safeguard. We parse all foreign key relationships in the database schema $\mathcal{D}$. For any pair of tables $(T_{i},T_{j})$ present in $\mathcal{D}_{union}$, we automatically add the corresponding primary and foreign key columns that link them. This step transforms the set of isolated tables into a connected graph, providing a solid topological foundation for the subsequent SQL generation phase. This final, closed schema is defined as:
@@ 3.3. N-version Programming for SQL Generation
ICL-based SQL Generation. This generator leverages in-context learning (ICL), analogous to case-based reasoning (Kolodner, 1993). By providing the LLM with relevant examples, we ground its generation in proven patterns. To eliminate domain-specific bias, we adopt a masked retrieval strategy following DAIL-SQL (Gao et al., 2024). Specifically, we mask table and column names in the user question with a special token (i.e., <MASK>) to focus solely on structural logic. We then encode these masked questions using the all-MiniLM-L6-v2 (Reimers and Gurevych, 2019) model and retrieve the top-$K$ most similar examples from the training set of the corresponding benchmark (e.g., BIRD-Train) to serve as few-shot demonstrations. This is formalized as:
@@ 3.4. SQL Unit Testing and Revision via Tool-Chain
LLM-generated SQL often contains critical errors (e.g., incorrect JOINs), yet LLMs struggle to self-correct due to inherent confirmation bias (Liu et al., 2025b; Qu et al., 2025; Zhang et al., 2025b). Analogous to the principle that code requires independent validation, we introduce SQL Unit Testing and Revision via Tool-Chain. Drawing on Unit Testing (Runeson, 2006), we treat functional SQL components (e.g., JOIN logic, syntax) as testable “units”. The Tool-Chain externalizes verification through a deterministic chain of specialized checkers, functioning as an automated test suite to systematically detect errors and guide targeted revisions.
@@
Table 2. Tool-chain of specialized checkers organized by verification stage.
@@
Stage 1: Syntax & Execution (Fail-Fast)
Syntax Checker Syntax validity and execution errors SELECT col, COUNT(*) (missing GROUP BY); SELECT * FORM table (typo)
Stage 2: Logical Verification
JOIN Checker Non-standard or complex joins ON T1.id = T2.id OR ...; ON col IN (SELECT ...)
ORDER-BY Checker Logical conflicts in sorting and limits ORDER BY COUNT(*) LIMIT 1 (validate aggregation-based sorting)
Time Checker Invalid temporal function usage STRFTIME vs. DATETIME; invalid date-format comparisons
Stage 3: Quality & Robustness
SELECT Checker Ambiguity removal SELECT * $\rightarrow$ SELECT col_a, col_b
MaxMin Checker Efficiency optimization WHERE col = (SELECT MAX(...)) $\rightarrow$ ORDER BY col DESC LIMIT 1
NULL Checker Missing NULL guards Add WHERE col IS NOT NULL to sorting keys
Result Checker Empty-result verification Queries returning zero rows (triggering re-evaluation of constraints)
Our approach centers on a suite of specialized, deterministic programs we term “Checker Tools”. These checkers are not applied randomly; they are orchestrated into a sequential “Syntax $\to$ Logic $\to$ Quality” cascade. This order is critical for a “fail-fast” verification: it ensures that fatal execution errors are resolved first before the system attempts to diagnose more subtle logical or stylistic discrepancies. The complete, ordered sequence is detailed in Table 2. The chain begins with the most fundamental validation: the Syntax Checker ensures the query is syntactically valid and executable. Notably, this stage also acts as an implicit guard for aggregation errors (e.g., missing GROUP BY), as these typically trigger execution failures that are immediately captured and repaired. Once executable, the chain proceeds to Logical Verification. The JOIN Checker flags non-standard conditions, while the ORDER-BY Checker validates sorting logic and corrects invalid patterns (e.g., misaligned LIMIT clauses). Finally, the chain addresses Quality and Robustness issues. The SELECT Checker replaces ambiguous wildcards like SELECT * with specific column names. Concurrently, the NULL Checker and Result Checker inject guards for potential NULL issues and flag queries that produce empty results, ensuring the final output is not only correct but also practically useful.
@@
    The LLM is then invoked in a special “revision mode”. It receives the original context, the faulty SQL $\mathcal{S}_{cand}$, and the explicit directive $d_{err}$ from the checker. The LLM’s task is not to find the error, but to fix it based on the directive. It is formalized as:
@@ 3.5. Confidence-aware SQL Selection
The preceding phases of our framework produce a set of high-quality, revised SQL candidates, $\mathcal{C}_{revised}$. However, these candidates may not be identical and could yield different execution results. The most common approach for selecting a final query is self-consistency (Xie et al., 2025a; Li et al., 2025a; Gao et al., 2024), where all candidates are executed, and the query corresponding to the most frequent result is chosen. This method, while a strong baseline, has a critical flaw: the most popular answer is not always the correct one, especially for complex problems where multiple generation paths might converge on the same plausible but incorrect logic. This final challenge mirrors the release stage in a software lifecycle, which is governed by a Quality Gate (Schnell and Reinhart, 2016). A quality gate’s purpose is not merely to accept the most-voted-for version, but to enforce a set of objective quality criteria before a product is released. Similarly, our Confidence-aware SQL Selection phase acts as this quality gate for the generated SQL. It overcomes the flaws of simple majority voting by using the initial vote’s confidence as a quality metric to guide a more reliable, adaptive selection process.
@@
(1) Execution-Guided Prompting. Instead of a neutral comparison, we construct an asymmetric prompt guided by execution consensus. We identify the top-$K$ execution clusters and randomly select one representative SQL from each cluster to form the candidate set. The prompt explicitly injects the execution confidence as context (e.g., informing the LLM that Candidate A has higher execution confidence than Candidate B). We instruct the LLM to treat this high-confidence consensus as a prior, trusting it unless specific logical evidence—such as a contradiction between the generated SQL and the user’s natural language requirement—demonstrably proves it factually incorrect.
(2) Perform Pairwise Adjudication. Next, the LLM performs pairwise comparisons for the top-$K$ candidates. To ensure reliability and mitigate the inherent stochasticity of LLM outputs, we employ a self-consistency mechanism. For each pair $(\mathcal{S}_{i},\mathcal{S}_{j})$, we sample three independent judgments and define the final stable vote, $V(\mathcal{S}_{i},\mathcal{S}_{j})$, as the majority outcome. The vote $V(\mathcal{S}_{i},\mathcal{S}_{j})$ yields 1 if $\mathcal{S}_{i}$ is superior and 0 if $\mathcal{S}_{j}$ is superior. From these consistent pairwise results, we compute an aggregate win rate for each candidate $\mathcal{S}_{i}$ as its average score against all other competitors in the top-$K$ set:
@@ 4. Experiments
This section evaluates the performance and reliability of DeepEye-SQL. We first detail the experimental setup, followed by a comparative analysis against state-of-the-art methods. Finally, we conduct ablation studies and efficiency analyses to validate the contribution of individual components.
@@ 4.1. Experimental Setup
Datasets. We evaluate the performance of DeepEye-SQL on three challenging Text-to-SQL benchmarks: BIRD (Li et al., 2024c), Spider (Yu et al., 2018), and the newly released Spider 2.0 (Lei et al., 2025). BIRD is a large-scale benchmark designed to mirror complex, real-world scenarios. It contains 12,751 unique question-SQL pairs across 95 databases from over 37 professional domains. Its databases are notably large and feature messy data and intricate schemas, making it a difficult test for grounding and robustness. Spider is a foundational large-scale, cross-domain benchmark in the field. It consists of 10,181 questions and 5,693 unique, complex SQL queries across 200 databases covering 138 different domains. Spider 2.0 represents a significant leap towards enterprise-grade challenges. Unlike its predecessor, it introduces complex schema and diverse dialects (e.g., BigQuery, Snowflake) to simulate real-world data warehousing environments. It is divided into two subsets: Lite and Snow. Following prior works (Pourreza et al., 2025; Liu et al., 2025c; Cao et al., 2024; Deng et al., 2025), we use the development set of BIRD (BIRD-Dev), the test set of Spider (Spider-Test), and both the Snow and Lite subsets of Spider 2.0 for our main evaluation.
@@
Evaluation Metrics. Following prior works (Pourreza et al., 2025; Liu et al., 2025c; Cao et al., 2024), our primary evaluation metric is Execution Accuracy (EX). For BIRD and Spider, a generated SQL is considered correct if its execution result is strictly equivalent to the ground truth. For Spider 2.0, we adhere to its official evaluation protocol (Lei et al., 2025), which employs a relaxed comparison standard that permits redundant columns in the result set, acknowledging the practical needs of data warehousing environments. To measure the potential of our N-version Programming for SQL Generation module, we report the Upper-bound EX, which is the execution accuracy an oracle would achieve by always selecting the correct SQL from the generated candidates (Pourreza et al., 2025). For evaluating the Robust Schema Linking module, we use Table/Column Recall, the proportion of ground-truth tables and columns correctly identified. Finally, to assess practical efficiency, we measure Token Cost, corresponding to the number of tokens processed by LLMs.
@@
Baselines. We compare DeepEye-SQL against SOTA baselines from two paradigms (Table 3). The first is fine-tuning-based methods, which are trained on in-domain data, including strong competitors like XiYan-SQL (Liu et al., 2025c), CHASE-SQL (Pourreza et al., 2025), and OmniSQL (Li et al., 2025b). The second is prompting-based methods, which, like DeepEye-SQL, are training-free. This group includes methods like Alpha-SQL (Li et al., 2025a), RSL-SQL (Cao et al., 2024), and CHESS (Talaei et al., 2024). Furthermore, for the Spider 2.0 evaluation, we compare against specialized agentic frameworks designed for enterprise-grade complexity, including Spider-Agent (Lei et al., 2025), ReFoRCE (Deng et al., 2025) and LinkAlign (Wang and Liu, 2025).
@@
Implementation Details. All experiments are conducted on an Ubuntu 22.04.3 LTS server with 512GB of RAM and dual 40-core Intel(R) Xeon(R) Platinum 8383C CPUs. We deploy all open-source LLMs locally using the vllm (Kwon et al., 2023) framework on 8 NVIDIA A100 GPUs, each with 80GB of memory, and accelerate inference using a tensor parallelism of 8. To validate the robustness and generalizability of our framework, we integrated DeepEye-SQL with three distinct models from two leading series: Gemma3-27B-Instruct (Kamath et al., 2025), Qwen2.5-Coder-32B-Instruct (Hui et al., 2024), and Qwen3-Coder-30B-A3B-Instruct (Yang et al., 2025). Additionally, for the Spider 2.0 benchmark, which requires extreme reasoning capabilities to navigate complex enterprise schemas, we adopted the DeepSeek-R1 model (DeepSeek-AI, 2025), consistent with the configuration of comparable baselines. Unless otherwise specified, the pipeline is configured as follows. For Semantic Value Retrieval, where we retrieve up to the top-5 most similar values for each TEXT-type column, we use the Qwen3-Embedding-0.6B (Zhang et al., 2025c) model, and the similarity threshold ($\theta_{val}$) for Value-based Schema Linking is set to 0.98. The draft SQL for Reversed Schema Linking is produced by the ICL-based generator. For each LLM sub-task (e.g., Direct Schema Linking), we set the sampling budget to 8 with a sampling temperature of 0.7. Finally, in the Confidence-aware SQL Selection phase, which adjudicates between the top-2 ranked queries in low-confidence scenarios, the confidence shortcut threshold ($\theta_{conf}$) is set to 0.6.